wat do trace and rank of a matrix mean?????

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Unnati,

 Trace : Sum of the elements of A (square matrix of order n ) lying along the principle diagonal is called trace of A.

Ex -   matrix =  1    0     so trace = 1+5 =6

                       0   5

Rank : No. r is said to be the rank of a m*n matrix A if,

1 every square sub matrix of order (r+1) or more is singular and

2 there exists at least one square submatrix of order r which is non singular

or rank of a matrix is the order of highest order non singular sub matrix

Ex - A = 1    2    3

            4    5    6

            3    4    5

we have det. A =0 so, r(A) is less then 3, also we have   5    6    a non singular square sub matrix of order 2 .

                                                                               4    5

hence rank of A = 2

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

 

 

SAGAR SINGH - IIT DELHI
13 years ago

Dear unnati,

Let ${\bf A}$ be an $n \times n$ square matrix:

\begin{displaymath}{\bf A}=\left[ \begin{array}{cccc} a_{11} & a_{12} & \cdot &... ...n1} & a_{n2} & \cdot & a_{nn} \end{array} \right]_{n\times n} \end{displaymath}


where

\begin{displaymath}\left[ \begin{array}{c} a_{1j} a_{2j} ... a_{nj} \end{array}\right] \end{displaymath}


is the jth column vector and

\begin{displaymath}[a_{i1}\;a_{i2}\;\cdots\;a_{in}]\end{displaymath}


is the ith row vector.

The n rows span the row space of ${\bf A}$ and the n columns span the column space of ${\bf A}$. The dimensions of these two spaces are the same and called the rank of ${\bf A}$:

\begin{displaymath}R=rank({\bf A}) \leq N \end{displaymath}

 

The determinant of $A$ is denoted by

\begin{displaymath}det({\bf A}) = \vert {\bf A} \vert \end{displaymath}


and we have

\begin{displaymath}\vert {\bf A}{\bf B} \vert=\vert{\bf A}\vert\; \vert{\bf B}\vert \end{displaymath}

 

$rank({\bf A}) < N$ if and only if

$det({\bf A})=0$.

The trace of $A$ is defined as the sum of its diagonal elements:

\begin{displaymath}tr({\bf A}) = \sum_{i=1}^n a_{ii} \end{displaymath}

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi